OpenStudy (anonymous):

A hoop rolls down an inclined plane. What fraction of its total KE is associated with its rotation? I=mr^2

OpenStudy (anonymous):

for a hoop k.Erot = 1/2 mv^2 u know how ???

OpenStudy (anonymous):

i don't

OpenStudy (anonymous):

K.E =1/2 Iw^2 I for hoop is mr^2 So K.E = mr^2w^2 r^2w^2 = v^2 K.E = 1/2 mv^2 w = omega Angular velocity I = moment of Inertia